OpenStudy (anonymous):

What value of x makes the equation true? two-eighths equals twelve over x ninety-six forty-eight eighteen fourteen Question 2 (Multiple Ch

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

2/8 = 12/x?

OpenStudy (mathmath333):

28=12/x x=3/7

OpenStudy (anonymous):

Ugm, I thought it was 48?

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

2/8 = .25 and 12/48 = .25...

OpenStudy (anonymous):

x = 48, you can check it to make sure.

OpenStudy (anonymous):

what so what is it

OpenStudy (anonymous):

x = 48, if you plugged in 48 for x you can simplfy 12/48 to 2/8, thus making it true.

OpenStudy (mathmath333):

u can write question in equation format

OpenStudy (mathmath333):

i am unable to understand ur queation

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

thanks for trying